OpenStudy (anonymous):

Trying to solve an identity with the sum and difference formula, (working one side): sin (pi/6) + x)) + sin (pi / 6) -x)) = cos x

OpenStudy (anonymous):

so we can start expanding out the formula we get ... sin(pi/6)cos(x) + sin(x)cos(pi/6) + sin(pi/6)cos(x) - cos(pi/6)sin(x) so now we get two terms because the other two terms cancel out. the term that cancels out is sin(x)cos(pi/6) so we simplify to sin(pi/6)cos(x) + sin(pi/6)cos(x) so ... 2sin(pi/6)cos(x) now we know that sin(pi/6) = 1/2 so 2* ( 1/2) =1 hence 1*cos(x) = cos(x)

OpenStudy (anonymous):

Thank you so much! I really appreciate your expertise! :)

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

:)